# Learning local equivariant representations for quantum operators

### v2

@@ III.3.5 Edge updates
Table 1: MAE (in meV) for Hamiltonian matrix predictions using SLEM and other methods on materials with DFT-LCAO basis up to $d$, $f$, and $g$ orbitals. Numbers in parentheses indicate parameter count. Data for HfO2 in DeepH-E3 and for both GaN and HfO2 in HamGNN are absent due to out-of-memory errors.
@@
(0.7M) (4.5M) ( 1.0 M) ( 4.5M) (8.3 M)
MoS2 0.34 0.14 0.46 0.55 0.63
Graphene 0.26 0.14 0.40 0.28 0.40
Si(300K) 0.10 0.07 0.16 0.10 0.40 (25.5 M)
@@
Table 2: MAE for predicting density matrix using SLEM on materials with DFT-LCAO basis up to $d$, $f$, and $g$ orbitals. Model settings align with those in Table LABEL:hamiltonian.
@@ IV.1 Benchmark the accuracy and data-efficiency
As shown in Table LABEL:hamiltonian, our SLEM model achieves state-of-the-art accuracy in Hamiltonian prediction, exhibiting the lowest mean absolute error (MAE) across all test systems compared to referenced methods. Notably, high accuracy is achieved with a relatively small model size (only 0.7 million (M) trainable parameters). Fig. 3 illustrates the band structure of silicon structures computed directly from the SLEM-predicted Hamiltonian, where the eigenvalues are indistinguishable from those obtained using DFT. For the density matrix, fitting results are presented in Table LABEL:density. The results demonstrate very high accuracy (order of 1e-5), approaching the machine precision limit of float32 numbers. In Fig. 3, we use the trained model to predict the density matrix and visualize its real-space distribution. This capability is particularly important for applications such as charge distribution analysis or tracking electron transfer via methods like differential charge density. Furthermore, the predicted density can be directly used for non-self-consistent field (NSCF) DFT calculations. The resulting band structure for silicon, as an example, is highly accurate and matches the DFT output, with a MAE of only 1.09 meV in eigenvalues compared to self-consistent DFT results. The overlap matrix, represented by invariant parameters in our SLEM model, achieves exceptionally high accuracy as demonstrated in Table LABEL:overlap, approaching the machine precision limit for float32 numbers. Notably, our simplified parameterization enables this high accuracy with only a minimal increase in model complexity. For instance, in a silicon model designed to fit only the Hamiltonian, our typical parameter count is 0.7 M. The inclusion of overlap matrix prediction adds merely 0.01 M parameters, representing an increase of only about 1.4% in total model size.
Table 3: MAE for predicting overlap matrix using SLEM’s parameterization on materials with DFT-LCAO basis up to $d$, $f$, and $g$ orbitals. Model settings align with those in Table LABEL:hamiltonian.
@@
Table 4: Comparison of validation MAE (in meV) for SLEM model and literature methods trained on randomly split datasets [11] with varying training ratios. Model settings align with those in Table LABEL:hamiltonian.
@@
HamGNN 0.41 0.32 0.37 0.42 0.60
Additionally, due to the strict localization scheme, our SLEM model demonstrates superior data efficiency. Specifically, training the SLEM model requires fewer data points generated via DFT calculations. To quantify the data-efficiency performance, we designed an experiment by randomly splitting the dataset into subsets containing 20%, 40%, 60%, and 80% of the original training data. We then trained the strictly local SLEM model and other methods using these subsets and validated their performance on the same validation sets. The results, presented in Table LABEL:data-efficiency, illustrate the remarkable data-efficiency of our method. This indicates that using SLEM allows users to generate a smaller, cost-effective training set. Furthermore, the excellent data efficiency and transferability of SLEM make it particularly suitable as a backbone model for developing a universal DFT model, especially for systems involving heavy elements.
@@ IV.2 Efficiency and Scalability
Firstly, for efficiency, the implementation of SO(2) convolution reduces the tensor product computational complexity from $O(L^{6})$ to $O(L^{3})$, which is then further reduced by the parallelization of matrix operations to near $O(L)$ benefiting from PyTorch. In Fig. 4, we compare the wall time and GPU memory consumed by tensor product operations using SO(2) convolution and the normal method employed in DeepH-E3 [8] and E3NN [42]. Clearly, tensor production via SO(2) convolution is far more efficient than conventional methods. This reduction of cost makes our method very efficient, which is capable of handling all possible basis choices in LCAO DFT. We also record the memory and time required when training on some of the typical systems. The results are displayed in Figure.5. We can see, that in all systems, our model trains faster in time while requiring less memory. Notably, the larger the basis set, the greater the advantage our method has in both memory and time efficiency.
Figure 5: Comparison of training time per iteration and memory consumption among SLEM, DeepH-E3 [8], and HamGNN [10]. For GaN, HamGNN data is absent due to out-of-memory error.
@@
Therefore, it is possible to divide the whole atomic system into sub-graphs and compute their node and edge features independently on different devices. This is extremely important when expanding the simulation power of DFT to very large systems. In practice, we find that for HfO2 with $4s2p2d2f1g$ basis, a typical model of 1.7M parameters can only predict the quantum tensors of structures up to $10^{3}$ atoms on devices with 32GB memory. Despite the linear dependency, the inference on a system with $10^{4}\sim 10^{7}$ would require a memory with potentially over 300 GB, which is only available when parallelised to multiple GPUs. Therefore, a strictly localized model such as SLEM holds significant potential for expanding simulation system sizes.

### v3

@@ III.3.5 Edge updates
Table 1: MAE (in meV) for Hamiltonian matrix predictions using SLEM and other methods on materials with LCAO basis up to $d$ orbitals. Numbers in parentheses indicate parameter count.
@@
(0.7M) (4.5M) ( 1.0 M) ( 4.5M) (2.8M)
MoS2 0.34 0.14 0.46 0.55 1.20
Graphene 0.26 0.14 0.40 0.28 0.35
Si(300K) 0.10 0.07 0.16 0.10 0.19
Table 2: MAE (in meV) for Hamiltonian matrix predictions using SLEM and DeepH-E3 models [8] on materials with LCAO basis up to $f$ and $g$ orbitals. Numbers in parentheses indicate parameter count. The MAE of HfO2 in DeepH-E3 is absent due to out-of-memory errors.
@@
Table 3: MAE for predicting density matrix using SLEM on materials with LCAO basis up to $d$, $f$, and $g$ orbitals. Model settings align with those in Table LABEL:hamiltonian.
@@ IV.1 Benchmark the accuracy and data-efficiency
Table 4: MAE for predicting overlap matrix using SLEM’s parameterization on materials with LCAO basis up to $d$, $f$, and $g$ orbitals. Model settings align with those in Table LABEL:hamiltonian.
@@
Table LABEL:hamiltonian presents a comparison of mean absolute error (MAE) values in Hamiltonian prediction for graphene, MoS2, and Si systems, whose LCAO basis extends up to $d$ orbitals, among the SLEM, DeepH-E3 [8], and HamGNN [10] methods. Our SLEM model achieves state-of-the-art accuracy, exhibiting the lowest MAE across all systems. Notably, this high accuracy is achieved with a relatively small model size of only 0.7 million (M) trainable parameters. Furthermore, we extended our comparison between the SLEM and DeepH-E3 methods to include GaN and HfO2 systems, where the LCAO basis extends up to $f$ and $g$ orbitals, respectively. As shown in Table LABEL:hamiltonian_fg, the SLEM model consistently presents the lowest MAE, further demonstrating its high accuracy and versatility across different orbital complexities. Fig. 3 illustrates the band structure of silicon structures computed directly from the SLEM-predicted Hamiltonian, where the eigenvalues are indistinguishable from those obtained using DFT. For the density matrix, fitting results are presented in Table LABEL:density. The results demonstrate very high accuracy (order of 1e-5), approaching the machine precision limit of float32 numbers. In Fig. 3, we use the trained model to predict the density matrix and visualize its real-space distribution. This capability is particularly important for applications such as charge distribution analysis or tracking electron transfer via methods like differential charge density. Furthermore, the predicted density can be directly used for non-self-consistent field (NSCF) DFT calculations. The resulting band structure for silicon, as an example, is highly accurate and matches the DFT output, with a MAE of only 1.09 meV in eigenvalues compared to self-consistent DFT results. The overlap matrix, represented by invariant parameters in our SLEM model, achieves exceptionally high accuracy as demonstrated in Table LABEL:overlap, approaching the machine precision limit for float32 numbers. Notably, our simplified parameterization enables this high accuracy with only a minimal increase in model complexity. For instance, in a silicon model designed to fit only the Hamiltonian, our typical parameter count is 0.7 M. The inclusion of overlap matrix prediction adds merely 0.01 M parameters, representing an increase of only about 1.4% in total model size.
Table 5: Comparison of validation MAE (in meV) for SLEM model and DeepH-E3 [8] method trained on randomly split datasets [11] with varying training ratios. Model settings align with those in Table LABEL:hamiltonian.
@@
Additionally, the strict localization scheme of our SLEM model confers superior data efficiency, requiring fewer DFT-calculated data points for training. To quantify this efficiency, we conducted an experiment using randomly split subsets of the original training data, comprising 20%, 40%, 60%, and 80% of the full dataset. We trained both the SLEM model and DeepH-E3 method on these subsets and evaluated their performance on consistent validation sets. Table LABEL:data-efficiency presents the results, demonstrating SLEM’s maintained high accuracy across all subsets, thus highlighting its remarkable data efficiency. This data efficiency implies that SLEM users can generate smaller, more cost-effective training sets. Moreover, SLEM’s excellent data efficiency and transferability make it particularly well-suited as a backbone for developing universal DFT models, especially for systems involving heavy elements.
@@ IV.2 Efficiency and Scalability
Firstly, for efficiency, the implementation of SO(2) convolution reduces the tensor product computational complexity from $O(L^{6})$ to $O(L^{3})$, which is then further reduced by the parallelization of matrix operations to near $O(L)$ benefiting from PyTorch. In Fig. 4, we compare the wall time and GPU memory consumed by tensor product operations using SO(2) convolution and the normal method employed in DeepH-E3 [8] and E3NN [42]. Clearly, tensor production via SO(2) convolution is far more efficient than conventional methods. This reduction of cost makes our method very efficient, which is capable of handling all possible basis choices in LCAO DFT. We also record the memory and time required when training on some of the typical systems and compare with that of DeepH-E3 [8]. The results are displayed in Fig. 5. We can see, that in all systems, our model trains faster in time while requiring less memory. Notably, the larger the basis set, the greater the advantage our method has in both memory and time efficiency.
Figure 5: Comparison of training time per iteration and memory consumption with SLEM and DeepH-E3  [8] models.
@@
Therefore, it is possible to divide the whole atomic system into sub-graphs and compute their node and edge features independently on different devices. This is extremely important when expanding the simulation power of DFT to very large systems. In practice, we find that for HfO2 with $4s2p2d2f1g$ basis, a typical model of 1.7M parameters can only predict the quantum operators of structures up to $10^{3}$ atoms on devices with 32GB memory. Despite the linear dependency, the inference on a system with $10^{4}\sim 10^{7}$ would require a memory with potentially over 300 GB, which is only available when parallelised to multiple GPUs. Therefore, a strictly localized model such as SLEM holds significant potential for expanding simulation system sizes.